OpenStudy (anonymous):

How would i go about finding cos of 7pi/4 ? (And for sin to but if its the same process then cos is only needed and i'll connect the dots)

OpenStudy (anonymous):

look at the last page of the attached cheat sheet. find \(\frac{7\pi}{4}\) on the unit circle the first coordinate is cosine, the second coordinate is sine

OpenStudy (anonymous):

or you have to memorize a triangle and put it in the unit circle in the appropriate place 

OpenStudy (anonymous):

right \(\frac{\sqrt{2}}{2}\) and down \(\frac{\sqrt2}{2}\) so \[\cos(\frac{7\pi}{4})=-\frac{\sqrt2}{2}\] and \[\sin(\frac{7\pi}{4})=\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

Oh wow thanks, i feel dumb for not looking over the entire unit circle, for some reason i was thinking 7pi/4 was not on it.

OpenStudy (anonymous):

all numbers are on it sometimes you just have to keep going round and around

OpenStudy (anonymous):

One thing, would -\[-\sin7\pi/4 + 7\cos7\pi/4= 7\sqrt{2}/2 ?\]

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

we have the numbers we need you will get \[\frac{\sqrt2}{2}+7\frac{\sqrt2}{2}=\frac{8\sqrt2}{2}=4\sqrt2\]

OpenStudy (anonymous):

Oh thanks, i just realized this when looking at an example on the math program? didn't realize an invisible one (That is what it is right?)